# Towards Effective Prompt Stealing Attack against Text-to-Image Diffusion Models

### v1

@@ V-C Ablation Study
Impact of Dynamic Modifiers. Figure 8 and 9 show the impact of the number of captions that are used to generate dynamic modifiers, ranging from 0 to 400, on Prometheus’s performance with three metrics. We have two observations. First, including dynamic modifiers boosts the stealing performance. As shown in all cases, when adopting dynamic modifiers (quantity = 100), the similarities in the three metrics outperform the method that excludes dynamic modifiers (quantity = 0). The reason is that our proposed dynamic modifiers enhance the prompt’s detail and positional information, providing a better description of the showcase. Second, Prometheus shows a stable performance trend as the modifier quantity increases from 100 to 400. The potential reason is that Prometheus extracts high-quality dynamic modifiers from the captions of the showcase, requiring only a limited number of captions. This reveals the high efficiency of Prometheus.
@@
[…]
Impact of Specific Dynamic Modifiers. We further validate the effectiveness of specific components of dynamic modifiers, $i.e.$, subject detail and position information. Here, we consider Prometheus with only one kind of dynamic modifiers by excluding the other kind during modifier collection. As shown in Table VI, we have two observations. First, including one or both types of dynamic modifiers improves performance. Subject detail and position information both have gains in image appearance and prompt fidelity, and the gains are even greater when they are both included. Second, subject detail provides more gain than position details. The position information includes position details and azimuth information, where the former is covered by the subject detail to a certain degree.
@@
Impact of Contextual Matching. We replace our proposed contextual matching method in Prometheus with a vanilla approach that directly calculates cosine similarity between the modifiers and showcases [25]. Table VII presents the results across three metrics. As the matching method transitions from vanilla to our approach, the attack performance improves. Notably, our contextual matching outperforms the vanilla method across all proxy models. A key highlight is the significant increase in prompt semantic similarity for prompts stolen with FLUX, rising from 0.789 to 0.814. The core reason is that our contextual matching considers modifiers’ contextual semantics so that it improves the prompt similarity.
Impact of Proxy Feedback. We now illustrate how the Proxy feedback influences the performance of Prometheus. We consider two baselines to replace our greedy search, i.e., direct splicing and random sampling. Direct splicing implies that the attacker directly splices the top-$k$ modifiers with the highest correlation to the subject, while random sampling chooses $k$ modifiers at random. We set $k$ to 20 for these two baselines and the sampling pool of random sampling to the best 200 modifiers, in line with the query budget of Prometheus. Figures 10 and 11 show the comparison results on three metrics. As all the modifiers are refined with our local contextual matching, which is based on modifier-showcase cosine similarity, all the methods achieve a rather good performance. This can be reflected in Figure 12(c). Beyond that, we observe our proposed method significantly improves the image and prompt similarity, as shown in Figures 12(a) and 12(b). The reason is twofold. First, although our local ranking provides a highly correlated match (a correlation coefficient of 0.65), it is insufficient to directly adopt the top modifiers as the final choice, as there is still a non-negligible gap. More detailed analysis can be found in Section IV-D. Second, Prometheus remedies this issue by leveraging the feedback from Proxy. This helps keep the truly relevant modifiers while dropping the irrelevant ones, and thus improves the appearance alignment and prompt fidelity.
@@
Impact of Query Budget. Figures 12 and 13 illustrate the performance of Prometheus across all three metrics as the query budget increases from 50 to 300. We observe that when the query budget rises from 50 to 100, Prometheus demonstrates significant improvements in image and prompt similarities. However, the performance stabilizes as the budget increases from 100 to 300. This indicates the efficiency of Prometheus, as it requires relatively few queries to achieve satisfactory stealing results. This also validates the effectiveness of our local contextual matching, whose coarse-grained filtering restricts an exact range, which further facilitates the fine-grained filtering in Proxy query.
@@ V-D User Study
We conducted a user study to gain deeper insights into users’ preferences regarding the recovered prompts. The question asked of the interviewees is “How similar do you think image A is to image B?”. To mitigate potential bias, we did not inform them of our task or what these images represented. The preference is divided into five levels, which are specified in the Appendix. We conducted our investigation anonymously, and finally, we gathered 103 valid answer sheets. The detailed user information is depicted in A-B. Figure 14 shows the user study results. We can tell that our Prometheus achieves the highest similarity according to respondents’ preference, indicating its practicality in the real world. Take-away: With sufficient ethical consideration and under fair settings, Prometheus achieves the highest alignment with human perceptions.
@@ VI-A Conventional Defenses
We first consider some existing defenses as below. (1) Random noise. We generate a normally distributed noise with the same threshold of $\epsilon$ to ensure a similar visual disturbance. (2) Puzzle effect. This defense covers the showcase with the puzzle effect, which is applied in practice [4]. (3) Text watermark. This is also one defense method used in reality [21]. It covers the image with repetitive text indicating the ownership. We present the visual effect of these defenses in Figure 15. Note that we do not consider PromptShield proposed in PromptStealer [43], as it assumes the defender has white-box access to the exact modifier prediction model, which is beyond our threat model. Instead, we have a thorough analysis in A-C. We also exclude methods for prompt copyright protection (e.g., PromptCARE [52]), as they can only verify the ownership of prompts, but cannot prevent privacy leakage. Another method, Glaze [42], is a perturbation-based method to prevent unauthorized usage of images. It optimizes an invisible perturbation to mislead the showcase feature to a target style. Unfortunately, the code for Glaze is not open-sourced, so we do not include it in our evaluation.
@@ VI-B Dedicated Defense
Given the ineffectiveness of the above defenses, we propose a new solution dedicated to Prometheus, dubbed PromptGuard. It aims to preserve the utility of showcasing images for benign users while preventing the reconstruction of prompts by malicious users. The core insight of PromptGuard is to add an invisible perturbation $\epsilon$ to the showcase, which can disturb the feature in the embedding spaces of the caption models ($\mathcal{B}$) and modifier matching model ($\mathcal{C}$). A defender can utilize any captioning model and modifier matching model to disrupt the embeddings. Its optimization goal can be formulated as:
@@
Defenses ${CLIP}_{img}(\downarrow)$ $LPIPS(\uparrow)$ $SBERT(\downarrow)$
No Defense 0.901 0.625 0.799
Random Noise 0.887 (-0.014) 0.775 (+0.150) 0.795 (-0.004)
Puzzle 0.882 (-0.019) 0.659 (+0.034) 0.799 (-0.000)
Watermark 0.829 (-0.072) 0.735 (+0.110) 0.788 (-0.011)
PromptGuard 0.824 (-0.077) 0.664 (+0.039) 0.774 (-0.025)
$\text{Prometheus}^{*}$ 0.915 (+0.014) 0.597 (-0.028) 0.804 (+0.005)
Limitation of PromptGuard. The success of PromptGuard heavily relies on the assumption that the defender knows the caption models ($\mathcal{B}$) and modifier matching model ($\mathcal{C}$) used by the attacker. However, this assumption does not hold water as the attacker can try other models when he realizes that the chosen $\mathcal{B}$ and $\mathcal{C}$ are shielded. Unfortunately, in this case, the perturbation generated has low transferability to the different caption models and modifier matching models [9, 51]. This significantly restricts the practical value of PromptGuard. For example, PromptGuard generates the perturbation using BLIP with ViT-base, while the attacker can adopt BLIP with a different structure (e.g., ViT-large) for prompt stealing. The defense results for this setting are shown in Table VIII (denoted as $\text{Prometheus}^{*}$). We observe that with a different caption model, the stealing performance rises again, even outperforming the no-defense setting. This reveals the resistance of Prometheus to PromptGuard and the urgent need for more effective defenses. More discussions about such attacks can be found in Appendix A-C.
@@
VII Conclusion, Discussion, and Future Work
@@
The main limitation of Prometheus lies in its running time. Unlike prior works that rely on fussy data collection and pre-training, Prometheus introduces a lightweight online stealing framework featuring dynamic modifier generation and a Proxy-integrated search process that is both more adaptive and cost-efficient. On average, Prometheus takes about two minutes to steal a single prompt. This can be further optimized through parallel caption generation and more efficient search algorithms. Additionally, this paper proposes a search-based framework for effective prompt stealing, overcoming the overfitting limitations of the traditional training–prediction pipeline. Future work could explore alternative search strategies to further enhance attack performance.
VIII Ethics Considerations and Compliance with the Open Science Policy
VIII-A Ethics Considerations
@@
    Prompt Engineer. Prompt stealing attacks may directly harm the prompt engineer’s commercial interest. First, an attacker can avoid paying for the prompt but steal it. Second, after stealing, the attacker can reload it to prompt markets to violate the original seller’s IP further.
@@
    Prompt Market. Prompt stealing attacks may harm the prompt markets’ commercial interest. First, stealing prompt instead of buying would lead to no commissions earned for the prompt market. Second, if the attacker loads the stolen prompt to another platform, it would attract potential users away.
@@
    Artist and Society. Prompt stealing attacks can extract the style from an image. Therefore, if an artist posts his/her masterpiece, the attacker could extract a prompt, which describes precisely their work. Then the attacker can create infinite images that are similar to the artist’s works.
@@
    Balancing Risks and Benefits. We systematically assess risks and benefits across all stakeholders. To minimize the risks, we bought the researched prompts and obtained permission from their creators to use these prompts for research purposes.
  • 3)
@@
    Fairness. We considered fairness in the selection of research subjects, and equitable distribution of the burdens and benefits of research according to individual need, effort, societal contribution, and merit. For our user study, we conduct this experiment anonymously while considering sexual orientation, etc.
@@
    Equity. We considered how the burdens and benefits of publicizing newly discovered vulnerability balance out.
Respect for Law and Public Interest.
  • 1)
    Compliance. We engage in due diligence to identify laws, regulations, contracts, and other private agreements that apply to our research and respect these restrictions.
  • 2)
    Transparency and Accountability. We clearly communicated the purposes of our research and behaved responsibly. Our research methodology, ethical evaluations, data collected, and results generated are documented and made available responsibly in accordance with balancing risks and benefits.
VIII-B Compliance with the Open Science Policy
@@
Open Sharing of Code and Data. We will make all artifacts involved in our research, including datasets, test cases, scripts, and source code, publicly available on the GitHub platform to support academic exchange and technological advancement. The models used in our experiments (Stable Diffusion 3.5, FLUX, ShuttleDiffusion) are all open-source and can be accessed and downloaded from HuggingFace. Additionally, the dataset used in this study is the HF-Prompt dataset, an open and non-sensitive dataset available on the HuggingFace platform, ensuring transparency and openness.
Reproducibility and Replicability. All artifacts necessary for reproducing our research findings will be made public. We will also provide detailed experimental records and documentation on the GitHub platform, including information on the experimental environment setup, dependencies, and parameter settings, to enable researchers to replicate our experiments.
Appendix A Experiment
A-A VLMasExpert

### v2

@@ V-C Ablation Study
Impact of Dynamic Modifiers.
Figure 8 and 9 show the impact of the number of captions that are used to generate dynamic modifiers, ranging from 0 to 400, on Prometheus’s performance with three metrics. We have two observations. First, including dynamic modifiers boosts the stealing performance. As shown in all cases, when adopting dynamic modifiers (quantity = 100), the similarities in the three metrics outperform the method that excludes dynamic modifiers (quantity = 0). The reason is that our proposed dynamic modifiers enhance the prompt’s detail and positional information, providing a better description of the showcase. Second, Prometheus shows a stable performance trend as the modifier quantity increases from 100 to 400. The potential reason is that Prometheus extracts high-quality dynamic modifiers from the captions of the showcase, requiring only a limited number of captions. This reveals the high efficiency of Prometheus.
@@
0.882 0.674 0.796
0.897 0.659 0.810
0.891 0.665 0.807
0.901 0.653 0.814
Impact of Specific Dynamic Modifiers.
We further validate the effectiveness of specific components of dynamic modifiers, $i.e.$, subject detail and position information. Here, we consider Prometheus with only one kind of dynamic modifiers by excluding the other kind during modifier collection. As shown in Table VI, we have two observations. First, including one or both types of dynamic modifiers improves performance. Subject detail and position information both have gains in image appearance and prompt fidelity, and the gains are even greater when they are both included. Second, subject detail provides more gain than position details. The position information includes position details and azimuth information, where the former is covered by the subject detail to a certain degree.
@@
Impact of Contextual Matching.
We replace our proposed contextual matching method in Prometheus with a vanilla approach that directly calculates cosine similarity between the modifiers and showcases [47]. Table VII presents the results across three metrics. As the matching method transitions from vanilla to our approach, the attack performance improves. Notably, our contextual matching outperforms the vanilla method across all proxy models. A key highlight is the significant increase in prompt semantic similarity for prompts stolen with FLUX, rising from 0.789 to 0.814. The core reason is that our contextual matching considers modifiers’ contextual semantics so that it improves the prompt similarity.
@@
Impact of Proxy Feedback.
We now illustrate how the Proxy feedback influences the performance of Prometheus. We consider two baselines to replace our greedy search, i.e., direct splicing and random sampling. Direct splicing implies that the attacker directly splices the top-$k$ modifiers with the highest correlation to the subject, while random sampling chooses $k$ modifiers at random. We set $k$ to 20 for these two baselines and the sampling pool of random sampling to the best 200 modifiers, in line with the query budget of Prometheus. Figures 10 and 11 show the comparison results on three metrics. As all the modifiers are refined with our local contextual matching, which is based on modifier-showcase cosine similarity, all the methods achieve a rather good performance. This can be reflected in Figure 12(c). Beyond that, we observe our proposed method significantly improves the image and prompt similarity, as shown in Figures 12(a) and 12(b). The reason is twofold. First, although our local ranking provides a highly correlated match (a correlation coefficient of 0.65), it is insufficient to directly adopt the top modifiers as the final choice, as there is still a non-negligible gap. More detailed analysis can be found in Section IV-D. Second, Prometheus remedies this issue by leveraging the feedback from Proxy. This helps keep the truly relevant modifiers while dropping the irrelevant ones, and thus improves the appearance alignment and prompt fidelity.
Impact of Query Budget.
Figures 12 and 13 illustrate the performance of Prometheus across all three metrics as the query budget increases from 50 to 300. We observe that when the query budget rises from 50 to 100, Prometheus demonstrates significant improvements in image and prompt similarities. However, the performance stabilizes as the budget increases from 100 to 300. This indicates the efficiency of Prometheus, as it requires relatively few queries to achieve satisfactory stealing results. This also validates the effectiveness of our local contextual matching, whose coarse-grained filtering restricts an exact range, which further facilitates the fine-grained filtering in Proxy query.
@@ V-D User Study
We conducted a user study to gain deeper insights into users’ preferences regarding the recovered prompts. The question asked of the interviewees is “How similar do you think image A is to image B?”. To mitigate potential bias, we did not inform them of our task or what these images represented. The preference is divided into five levels, which are specified in the Appendix. We conducted our investigation anonymously, and finally, we gathered 103 valid answer sheets. The detailed user study setting, IRB exemption, and our effort to protect the volunteers are depicted in Appendix B-B. Figure 14 shows the user study results. We can tell that our Prometheus achieves the highest similarity according to respondents’ preference, indicating its practicality in the real world. Take-away: With sufficient ethical consideration and under fair settings, Prometheus achieves the highest alignment with human perceptions.
@@ VI-A Conventional Defenses
We first consider some existing defenses as below. (1) Random noise. We generate a normally distributed noise with the same threshold of $\epsilon$ to ensure a similar visual disturbance. (2) Puzzle effect. This defense covers the showcase with the puzzle effect, which is applied in practice [7]. (3) Text watermark. This is also one defense method used in reality [41]. It covers the image with repetitive text indicating the ownership. We present the visual effect of these defenses in Figure 15. We do not consider PromptShield proposed in PromptStealer [68], as it assumes that the attacker uses a modifier prediction model to steal the prompt and the defender has white-box access to the exact modifier prediction model, which does not apply to our attack and is beyond our threat model. Instead, we have a thorough analysis in B-C. We also exclude methods for prompt copyright protection (e.g., PromptCARE [83]), as they can only verify the ownership of prompts, but cannot prevent privacy leakage. Another method, Glaze [67], is a perturbation-based method to prevent unauthorized usage of images. It optimizes an invisible perturbation to mislead the showcase feature to a target style. Unfortunately, the code for Glaze is not open-sourced, so we do not include it in our evaluation.
@@ VI-B Dedicated Defense
Given the ineffectiveness of the above defenses, we propose a defense dedicated to Prometheus, dubbed PromptGuard. It aims to preserve the utility of showcasing images for benign users while preventing the reconstruction of prompts by malicious users. The core insight of PromptGuard is to add an invisible perturbation $\epsilon$ to the showcase, which can disturb the feature in the embedding spaces of the caption models ($\mathcal{B}$) and modifier matching model ($\mathcal{C}$). A defender can utilize any captioning model and modifier matching model to disrupt the embeddings. Its optimization goal can be formulated as:
@@
Defenses ${CLIP}_{img}(\downarrow)$ $LPIPS(\uparrow)$ $SBERT(\downarrow)$ $ASR(\downarrow)$
No Defense 0.901 0.625 0.799 62.5%
Random Noise 0.887 (-0.014) 0.775 (+0.150) 0.795 (-0.004) 58.3%(-4.2%)
Puzzle 0.882 (-0.019) 0.659 (+0.034) 0.799 (-0.000) 61.1%(-1.4%)
Watermark 0.829 (-0.072) 0.735 (+0.110) 0.788 (-0.011) 41.7%(-20.8%)
PromptGuard 0.824 (-0.077) 0.664 (+0.039) 0.774 (-0.025) 27.8%(-34.7%)
$\text{Prometheus}^{*}$ 0.915 (+0.014) 0.597 (-0.028) 0.804 (+0.005) 67.9%(+5.4%)
Limitation of PromptGuard.
The success of PromptGuard heavily relies on the assumption that the defender knows the caption models ($\mathcal{B}$) and modifier matching model ($\mathcal{C}$) used by the attacker. However, this assumption does not hold water as the attacker can try other models when he realizes that the chosen $\mathcal{B}$ and $\mathcal{C}$ are shielded. Unfortunately, in this case, the perturbation generated has low transferability to the different caption models and modifier matching models [17, 78]. This significantly restricts the practical value of PromptGuard. For example, PromptGuard generates the perturbation using BLIP with ViT-base, while the attacker can adopt BLIP with a different structure (e.g., ViT-large) for prompt stealing. The defense results for this setting are shown in Table VIII (denoted as $\text{Prometheus}^{*}$). We observe that with a different caption model, the stealing performance rises again, even outperforming the no-defense setting. This reveals the resistance of Prometheus to PromptGuard and the urgent need for more effective defenses. More discussions about such attacks can be found in Appendix B-C. We will discuss more effective defenses in future work.
@@
VII Potential Limitations and Future Directions
Time Complexity.
The primary limitation of Prometheus lies in its runtime. Unlike prior works that rely on fussy data collection and pre-training, Prometheus introduces a lightweight online stealing framework, featuring dynamic modifier generation and a Proxy-integrated search process that is both more adaptive and cost-efficient. On average, Prometheus takes about two minutes to steal a single prompt. Given that protected showcases are typically curated works, such as artistic prompts, which are relatively few compared to web-scale images, this time complexity would not become a practical obstruction for large-scale attacks. Moreover, as studied in Section V-C, our parameter choosing is redundant, so that the running time can be further optimized without noticeably affecting performance. Future work can focus on developing more efficient search strategies while preserving effectiveness.
Search Strategy.
This paper proposes a greedy search-based framework for effective prompt stealing, overcoming the overfitting limitations of the traditional training–prediction pipeline. We acknowledge that there may exist interdependence between modifiers so that they have a stronger combined effect than greedy search. However, we argue that, as our candidate modifiers are sampled offline with contextual matching before online search, which ensures every candidate already possesses a reasonably good descriptive capability for the showcase, the improvement of interdependent modifiers is likely to be marginal. However, doing so would significantly increase the computational cost. For example, in our setting with 200 candidates, even considering only pairwise combinations would theoretically slow down the search by approximately 100 times (i.e., 199/2). Given that the potential gain is limited, such a computational cost is unacceptable to a large extent. Future work could explore alternative search strategies beyond our adopted greedy search to further enhance attack performance while preserving effectiveness.
VIII Conclusion
@@
Acknowledgment
We thank all the reviewers for their constructive suggestions that helped improve this paper. We are also grateful to Boheng Li (Nanyang Technological University) for the insightful discussions held during the project. This research/project is supported by the National Research Foundation, Singapore, under its AI Singapore Programme (AISG Award No: AISG3-RPGV-2025-019 and No.AISG4-GC-2023-008-1B), and the Ministry of Education of Singapore under its Academic Research Fund Tier 2 (Grant No: MOE-T2EP20121-0008).
IX Ethics Considerations and Compliance with the Open Science Policy
IX-A Ethics Considerations
@@
    Prompt Engineer. Prompt-stealing attacks may directly harm the prompt engineer’s commercial interest. First, an attacker can avoid paying for the prompt but steal it. Second, after stealing, the attacker can reload it to prompt markets to violate the original seller’s IP further.
@@
    Prompt Market. Prompt-stealing attacks may harm the prompt markets’ commercial interest. First, stealing a prompt instead of buying would lead to no commissions earned for the prompt market. Second, if the attacker loads the stolen prompt to another platform, it would attract potential users away.
@@
    Artist and Society. Prompt-stealing attacks can extract the style from an image. Therefore, if an artist posts his/her masterpiece, the attacker could extract a prompt that describes their work precisely. Then the attacker can create infinite images that are similar to the artist’s works.
@@
    IRB Exemption. In our user study, all images used were non-explicit, stylized paintings representing everyday visual content and contained no inappropriate material (e.g., nudity or sexual themes). Accordingly, under standard ethical guidelines, this study qualified for exemption from formal IRB review, consistent with the “exempt review” category defined in U.S. IRB protocols (45 CFR 46) because it posed minimal risk to participants.
@@
    Volunteer Protection. To further safeguard participants, we implemented multiple protective measures. Before participation, all volunteers were informed that the study concerned human judgments of image similarity; the specific research purpose was deliberately described in general terms to prevent priming or potential misuse. Participants were briefed on the procedure and data handling, explicitly reminded that participation was voluntary and could be discontinued at any time, especially if they experienced any form of discomfort, and informed consent was obtained. No personally identifiable information was collected. Besides, all questionnaires were completed offline under direct supervision. Participants were allowed to withdraw immediately if they felt any discomfort, and their status was continuously monitored to ensure safety. Upon completion, participants retained the right to withdraw their responses, and we verified their physical and mental well-being before concluding the study.
IX-B Compliance with the Open Science Policy
@@
Open Sharing of Code and Data.
We will make all artifacts involved in our research, including datasets, test cases, scripts, and source code, publicly available on the GitHub platform to support academic exchange and technological advancement. The models used in our experiments (Stable Diffusion 3.5, FLUX, ShuttleDiffusion) are all open-source and can be accessed and downloaded from HuggingFace. Additionally, the dataset used in this study is the HF-Prompt dataset, an open and non-sensitive dataset available on the HuggingFace platform, ensuring transparency and openness.
Reproducibility and Replicability.
All artifacts necessary for reproducing our research findings will be made public. We will also provide detailed experimental records and documentation on the GitHub platform, including information on the experimental environment setup, dependencies, and parameter settings, to enable researchers to replicate our experiments.
Appendix A Method
A-A Problem Motivation
We specify the motivation for studying prompt stealing.
Prompts’ Value.
The value of well-crafted prompts can be validated from two perspectives. Firstly, there are numerous emerging prompt marketplaces and registered prompt engineers. The active prompt transactions prove the economic significance of exquisite prompts. Second, prompt engineering has become a recognized profession. For instance, Adobe offers prompt engineer positions with salaries between USD 162,000 and 301,200 [1], and similar listings appear on Indeed [6]. These trends highlight that well-crafted prompts are valuable intellectual assets.
Attack Scenario.
In some scenarios, users may have vague ideas of what they want to create but struggle to express them effectively, as well-designed prompts embed subtle linguistic cues that significantly influence output quality. Previous work [45] validates the “articulation barrier” when describing aspects like style or composition, especially for image generation. We target the scenario where users frequently encounter appealing AI-generated artworks online and may seek to reproduce their style, making prompt reconstruction both a realistic and attractive attack scenario.
A-B Privacy Protection Standpoint
From the privacy protection standpoint, prompt-stealing attack poses a significant threat to the confidentiality of prompts. Specifically, in many cases, a user’s prompts may interact with sensitive or private data during the generation process (e.g., rely on information retrieved from private databases or internal documents through in-context learning). Then, the recovered prompts include not only the ordinary user’s prompts but also the augmented sensitive information, as both are ultimately reflected in the generated images. For example, professional artists might use a prompt such as ’Generate a portrait in the style of my previous drafts (as shown in the provided examples)’. Stealing the prompt of such a generated portrait image could inadvertently expose the artist’s unique stylistic characteristics (i.e., the privacy in this case) embodied in those unreleased drafts.
A-C Dynamic Modifier Extraction
We detail the extraction process for dynamic modifiers in Algorithm 2. We first generate multiple descriptions of the showcase using a captioning model. We then apply the NLP toolkit Spacy [71] to each caption to obtain its part-of-speech (PoS) tags and dependency tree (DepTree). These two signals are subsequently used to extract subject details and position information, which together form the dynamic modifiers.
Algorithm 2 $\mathtt{Dynamic}\ \mathtt{Modifier}\ \mathtt{Extraction}$
1:caption model BLIP $\mathcal{B}$, showcase $s$, caption quantity $q$, and NLP analysis tool Spacy.
2:extracted dynamic modifier list $\mathcal{M}_{d}$.
[…]
6:for $p$ in $\mathcal{P}$ do
7:  $\mathtt{POS,DepTree}\leftarrow\mathtt{Spacy}(p)$
8:  $\mathcal{C}\leftarrow\mathtt{Spacy}.chunk(p)$ $\triangleright$ Get detail chunks
9:  $\mathcal{M}_{d}(d)\leftarrow\mathcal{M}_{d}(d)\cup\mathcal{C}$
10:  for $\mathtt{token}$ in $p$ do
11:   if $\mathtt{POS}(\mathtt{token})$ = “prep” then $\triangleright$ Find prepositions
12:     $\mathtt{child}\leftarrow\mathtt{DepTree}(\mathtt{token})$
13:     $c\leftarrow\mathtt{child}.join()$
14:     $\mathcal{M}_{d}(p)\leftarrow\mathcal{M}_{d}(p)\cup c$
15:   end if
16:  end for
17:end for
18:return $\mathcal{M}_{d}(d)$ and $\mathcal{M}_{d}(p)$
Appendix B Experiment
B-A VLMasExpert